OpenStudy (anonymous):

Maple syrup worth $6.00 a gallon and corn syrup worth .80 cents a gallon are used to make a mixture worth $2.36 a gallon. How many gallons of each kind of syrup are need to make 50 gallons of the mixture? {maple syrup: _______________________a0 gallons, corn syrup: _________________________a1gallons}

hero (hero):

m = gallons of maple syrup c = gallons of corn syrup Gallons of syrup m + c = 50 Cost of Syrup 6m + .8c = 2.36(m + c) 6m + .8c = 2.36(50) Solve this system: m + c = 50 6m + .8c = 2.36(50)

OpenStudy (anonymous):

M=14.2 C=35.8

hero (hero):

Close, but not quite

hero (hero):

The values for m and c are integers.

OpenStudy (anonymous):

can you explain?

hero (hero):

You need approximately 35 gallons of corn syrup and approximately 15 gallons of maple syrup to make 50 gallons of syrup.

hero (hero):

Actually....

hero (hero):

m + c = 50 6m + .8c = 50(2.36) 6m + 6c = 300 6m + .8c = 118 (6 - 6)m + (6 - .8)c = 300 - 118 0m + 5.2c = 182 5.2c = 182 c = 182/5.2 `c = 35` m + c = 50 m + 35 = 50 m = 50 - 35 `m = 15`